OpenStudy (marinedude42):

What is the slope of the line below? A. negative B. zero C. positive D. undefined

OpenStudy (marinedude42):

coordinates are (-2, 4) and (5,4 )

OpenStudy (compassionate):

Use the slope formula. 4 - 4/5-(-2) = 0/7 When you have a slope over 0, it's a zero slope.

OpenStudy (yttrium):

Just take note the slope of the line is the difference of y-coordinates divided by the difference of x-coordinates of that line. Mathematically, \[m = \frac{ y_{2} - y_{1} }{ x_{2}-x_{1} }\]

OpenStudy (mathmale):

If this is true: coordinates are (-2, 4) and (5,4 ) then \[m=\frac{ 4-4 }{x _{2}-x _{1} }\]Can you now finish the job of finding the slope of the line connecting your two points?

OpenStudy (marinedude42):

yea, got it

OpenStudy (mathmale):

Great. What is that slope?